# Infer-and-widen, or not?

### v2

@@ Abstract
In recent years, there has been substantial interest in the task of selective inference: inference on a parameter that is selected from the data. Many of the existing proposals fall into what we refer to as the infer-and-widen framework: they produce symmetric confidence intervals whose midpoints do not account for selection and therefore are biased; thus, the intervals must be wide enough to account for this bias. In this paper, we investigate infer-and-widen approaches in three vignettes: the winner’s curse, maximal contrasts, and inference after the lasso. In each of these examples, we show that a state-of-the-art infer-and-widen proposal leads to confidence intervals that are much wider than simple alternatives when all methods are tuned to yield identical randomized selection events. Furthermore, even an “oracle” infer-and-widen confidence interval — the narrowest possible interval that could be theoretically attained via infer-and-widen — is often wider than these alternatives.
@@ 1 Introduction
To formalize this, consider a random variable $Y$ with domain $\mathcal{Y}$, and a collection of parameters $\theta$ whose elements can be indexed by a (possibly infinite) index set $\mathcal{I}$. For a fixed index $i\in\mathcal{I}$, and for $\alpha\in(0,1)$, a classical confidence interval for $\theta_{i}$ is a confidence interval $\mathrm{CI}_{i}^{\alpha}(Y)$ with $1-\alpha$ coverage in the sense that
@@
Now, suppose that $i\in\mathcal{I}$ is no longer a fixed index; instead, some selection rule ${\mathcal{S}}:\mathcal{Y}\to\mathcal{I}$ uses the data to select a parameter on which to conduct inference. A confidence interval provides valid unconditional selective coverage (Fithian et al., 2017) at level $\alpha$ if
@@
Unconditional coverage is not guaranteed by classical confidence intervals. In this paper, we consider methods that yield unconditional guarantees, in the sense of 1.2. Some of these confidence intervals also yield (stronger) conditional selective guarantees, in the sense that
$\mathbb{P}\left(\theta_{i}\not\in\mathrm{CI}_{i}^{\alpha}(Y)~{}|~{}\mathcal{C}(Y)\right)\leq\alpha$ (1.3)
for an event $\mathcal{C}(Y)$ satisfying $\sigma\left(\mathcal{C}(Y)\right)\supseteq\sigma\left({\mathcal{S}}(Y)\right)$, where $\sigma(\cdot)$ denotes the sigma algebra generated by a random variable, i.e., the coverage guarantee in 1.3 conditions on at least the selection event; by the law of total expectation, conditional coverage 1.3 implies unconditional coverage 1.2.
@@ 1.1 The infer-and-widen framework
As presented in Box 1.1, the infer-and-widen framework involves (Step 1) defining a target of inference using some or all of the data, and then (Step 2) symmetrically adjusting the critical values arising from “classical” inference to account for the fact that the target of inference is not fixed in order to achieve unconditional coverage (1.2). The simplest examples of this strategy are the Bonferroni (Bonferroni, 1936), Holm-Bonferroni (Holm, 1979), and Šidák (Šidák, 1967) procedures, which are suitable when a finite number of targets of inference are considered (and in the case of Šidák, under additional assumptions). In the special case of inference on contrasts of a linear model, the Scheffé correction is available (Scheffé, 1959); this idea is further refined by the PoSI method (Berk et al., 2013). While conceptually attractive, these approaches tend to be quite conservative (Bonferroni, Holm-Bonferroni, Šidák), are limited to a linear model setting (Scheffé, PoSI), or are computationally intractable (PoSI).
Other recent infer-and-widen approaches have attempted to address these shortcomings. The local simultaneous inference approach of Zrnic and Fithian (2024b) improves upon the Bonferroni correction by focusing on the subset of hypotheses that are most likely given the observed data. By contrast, the algorithmic stability proposal of Zrnic and Jordan (2023) adapts to the details of the selection algorithm. By drawing connections with the field of differential privacy (Dwork and Roth, 2014), the “stability” of the approach for selecting the target of inference in Step 1 of Box 1.1 is quantified; this, in turn, leads to a stability-specific adjustment of the critical values in Step 2 of Box 1.1. In a sense, algorithmic stability achieves the lofty goals laid out in the PoSI proposal of Berk et al. (2013) by creating a framework to compute a selection-specific adjustment to the critical values (see also the online supplement).
[!htb] Box 1: Infer-and-widen. 1. A target of inference is selected from the data. 2. “Classical” inference is conducted on all of the data, with a symmetric adjustment of the critical values to account for the data-driven selection of the target of inference.
@@ 1.2 Other approaches for valid selective inference
Not all approaches that yield unconditional guarantees of the form (1.2) fall into the infer-and-widen framework. Some methods produce asymmetric confidence intervals around the biased point estimate: for instance, Benjamini et al. (2019) and Fuentes et al. (2018) construct an asymmetric Bonferroni correction that accounts for the direction of the bias. More recently, Zrnic and Fithian (2024a) extend the local simultaneous approach of Zrnic and Fithian (2024b) by additionally adapting the correction to the specific selection algorithm. The hybrid approaches of Andrews et al. (2019) and McCloskey (2024) apportion a fraction of the Type 1 error budget to rule out unlikely selection events, and then construct a conditionally-valid confidence interval on the likely hypotheses; note that this only guarantees unconditional coverage 1.2.
Furthermore, a number of methods yield coverage guarantees conditional on the selection event, i.e., (1.3). Some methods do this by obtaining independent training and test sets; then, the training set can be used to define the target of inference, and the test set to conduct inference on this target. Because the two sets are independent, we can conduct inference using the test set as though the target of inference were fixed, despite the fact that it is actually a function of the training data. Sample splitting, applicable in the setting of independent and identically distributed observations, provides one approach to obtain independent training and test sets (Cox, 1975). More recent data thinning proposals generalize the idea of sample splitting, and can be applied when only a single observation is available (Rasines and Young, 2023; Neufeld et al., 2024; Dharamshi et al., 2024).
In cases where it is either impossible or undesirable to create independent training and test sets, strategies that achieve conditional guarantees of the form 1.3 may still be available. For instance, a target of inference can be defined using all of the data, and then inference can be conducted conditional on the event that the data led to this particular target of inference. Examples include the proposals of Fithian et al. (2017); Lee et al. (2016); Chen and Witten (2023); Gao et al. (2024); Andrews et al. (2019). Alternatively, if a researcher is willing to define the target of inference using only a subset of the data or using a noisy version of the data, then data fission Leiner et al. (2023) or randomized conditional selective inference Tian and Taylor (2018); Panigrahi et al. (2024) may be applicable.
@@ 1.3 Contributions and organization
In this paper, we conduct an empirical and theoretical evaluation of infer-and-widen strategies. Since randomization is known to improve the performance of selective inference methods (Panigrahi et al., 2024; Tian and Taylor, 2018), we only consider randomized infer-and-widen approaches. The bottom line of our paper is as follows: when comparing infer-and-widen strategies to alternative selective inference methods that make use of identical selection events and yield conditional 1.3 (rather than unconditional 1.2 guarantees), the latter often outperform the former in terms of confidence interval width. Furthermore, the latter often outperform even the “oracle” infer-and-widen procedure, the narrowest possible interval that could be theoretically attained within the infer-and-widen framework.
The rest of this paper is organized as follows. In Section 2, we formalize the infer-and-widen framework. In Sections 3, 4, and 5, we compare its performance to alternative approaches in two vignettes considered in Zrnic and Jordan (2023), and in an example involving model selection via the lasso. In Section 6, we show analytically that the infer-and-widen interval can suffer from a highly biased midpoint. We close with a discussion in Section 7. Theoretical results are proven in the online supplement. Scripts to fully reproduce all numerical results can be found at https://github.com/rflperry/infer_and_widen.
@@ 2 Formalizing the infer-and-widen framework
We now formalize the infer-and-widen (I&W) framework, outlined in Box 1.1. This involves a modification to the classical confidence interval $\mathrm{CI}_{i}^{\alpha}$ in (1.1) to account for the fact that $\theta_{\mathcal{S}(Y)}$ is a data-driven parameter. Based on the particulars of the selection function $\mathcal{S}(\cdot):\mathcal{Y}\rightarrow\mathcal{I}$ and data $Y$, a modified $\alpha$-level, $H(\alpha,{\mathcal{S}},Y)$, is derived so that $\mathrm{CI}^{H(\alpha,{\mathcal{S}},Y)}_{\mathcal{S}(Y)}(Y)$ is unconditionally valid at level $\alpha$ in the sense of 1.2, i.e.,
@@
In a Bonferroni correction we have $H(\alpha,{\mathcal{S}},Y)=\alpha/|\mathcal{I}|$, and in the Holm-Bonferroni correction it is slightly larger. Since these approaches adapt neither to the selection rule $\mathcal{S}(\cdot):\mathcal{Y}\rightarrow\mathcal{I}$ nor to the data $Y$, the constant $H(\alpha,{\mathcal{S}},Y)$ is easy to compute but overly conservative.
The foundational PoSI proposal (Berk et al., 2013) considers the setting of linear regression with fixed regressors. They derive a constant $H(\alpha,{\mathcal{S}},Y)$ that enables inference on the coefficient(s) in any submodel, regardless of the form of $\mathcal{S}(\cdot)$ or observed data $Y$. However, for a particular selection rule $\mathcal{S}(\cdot)$, the resulting confidence interval may be quite conservative, and computing the constant $H(\alpha,{\mathcal{S}},Y)$ is challenging or infeasible.
Despite the notation $H(\alpha,{\mathcal{S}},Y)$ used in (2.1), the aforementioned infer-and-widen proposals result in confidence intervals whose widths do not depend on the particulars of the selection rule $\mathcal{S}(\cdot)$ nor data $Y$: that is, $H(\alpha,{\mathcal{S}},Y)$ could be written more simply as $H(\alpha)$.
In contrast are two newer approaches. The local simultaneous inference method of Zrnic and Fithian (2024b) apportions the Type-1 error budget $\alpha$, first to use the data $Y$ to rule out the most unlikely hypotheses, and then to perform a Bonferroni correction over a subset of all $\left|\mathcal{I}\right|$ hypotheses. We note that in this case, $H(\alpha,{\mathcal{S}},Y)$ could be written more simply as $H(\alpha,Y)$, as the correction does not take into account the selection rule. On the other hand, the algorithmic stability proposal of Zrnic and Jordan (2023) provides constants $H(\alpha,{\mathcal{S}})$ that are tailored to the particular selection rule ${\mathcal{S}}(\cdot)$ under consideration, although not to the data $Y$. The online supplement contains an overview of the algorithmic stability proposal; intuitively, a more “stable” selection rule ${\mathcal{S}}(\cdot)$ requires a less substantial adjustment to the critical values in Step 2 of Box 1.1.
@@ 3 Vignette #1: The winner’s curse
the selection rule is
$\mathcal{S}(Y):=\mathop{\mathrm{arg\,max}}_{i\in[n]}Y_{i},$ (3.2)
and our target of inference is $\mu_{\mathcal{S}(Y)}$. It is not hard to see that a $1-\alpha$ confidence interval for $\mu_{\mathcal{S}(Y)}$ that treats $\mathcal{S}(Y)$ as though it were fixed (rather than a function of the data) will fail to achieve the nominal coverage. This is the “winner’s curse” (Andrews et al., 2019).
Figure 1: Vignette #1 under (3.1) with $\mu=0$, $\sigma=1$, and the selection rule in 3.3. Results are averaged over (a) $250$ or (b, c) $1000$ simulated datasets. (a): We display the ratio of the average widths of the infer-and-widen 3.4 and data fission (DF) 3.7 intervals with $95\%$ coverage across values of the sample size $n$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the infer-and-widen interval is wider. Black regions represent pairs of $n$ and $c$ for which the infer-and-widen approach requires such a small Gaussian quantile that R returns an interval of infinite width. (b, c): Fixing $n=100$, we display the average width of the “oracle” infer-and-widen interval and the data fission interval 3.7 as a function of coverage. The 95% classical interval and the “attainable” 95% infer-and-widen interval based on algorithmic stability 3.4 are also displayed. The Laplacian noise variance is $0.33$ and $3$ in panels (b) and (c), respectively.
To achieve a “stable” algorithm, Zrnic and Jordan (2023) consider a randomized version of the selection rule (3.2) via the addition of centered Laplace noise,
@@
Proposition 1 (An infer-and-widen interval based on (3.3), given by (Zrnic and Jordan, 2023)).
@@
has $1-\alpha$ unconditional selective coverage for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ in the sense of 2.1.
We will now construct an alternative interval that makes use of exactly the same selection event, (3.3), using ideas from the data fission proposal of Leiner et al. (2023). Towards this end, note that $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)=\mathcal{S}(Y^{\mathrm{train}})$ where $Y^{\mathrm{train}}:=Y+\zeta$. Since $Y\mid Y^{\mathrm{train}}$ does not admit a tractable distribution, we will additionally condition on $\left({\rm sign}(\zeta_{1}),\ldots,{\rm sign}(\zeta_{n})\right)$. This yields a multivariate truncated normal distribution with independent marginals.
@@
has $1-\alpha$ conditional selective coverage for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ in the sense of 1.3 with $\mathcal{C}(Y)=\left(Y+\zeta,\delta\right).$
@@
Thus, both (3.4) and (3.7) provide confidence intervals for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ that attain $1-\alpha$ unconditional selective coverage. We should prefer the approach that yields narrower intervals. Figure 1(a) displays the ratio of the widths of the two intervals when $\mu=0$ and $\sigma^{2}=1$, as we vary the sample size $n$ and the Laplacian variance $2c^{2}$ in (3.3).
The infer-and-widen interval (3.4) involves tuning parameters $\nu$ and $\eta$; to construct Figure 1(a), for a fixed value of the noise $c$ and sample size $n$ in (3.3), we select the $\nu$ and $\eta$ that minimize the confidence interval width, subject to the constraints in Proposition 1. In the figure, black indicates values of $c$ and $n$ that necessitate a tiny quantile of the normal distribution, yielding numerical issues in R and therefore an interval (3.4) of apparently infinite width. The data fission intervals (3.7) are much narrower than the infer-and-widen intervals (3.4), across all values of $n$ and $c$ considered.
It is possible that the width of the infer-and-widen interval (3.4) is due to the fact that Zrnic and Jordan (2023) apply a series of inequalities in its derivation. Therefore, we now consider the “oracle” infer-and-widen interval: this is the narrowest possible interval centered at $Y_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ that achieves a given coverage. It is not computable in practice, as it requires knowledge of the true parameter value $\mu$. In Figures 1(b, c), we numerically compare the “oracle” infer-and-widen method to the data fission interval (3.7) under the selection rule (3.3) with variances $0.332c^{2}=$ 0.33 and $3$, respectively. The former method has a wider average width than the latter: thus, no infer-and-widen interval can outperform the data fission interval (3.4) in this setting.

### v3

@@ Abstract
In recent years, there has been substantial interest in the task of selective inference: inference on a parameter that is selected from the data. Many of the existing proposals fall into what we refer to as the infer-and-widen framework: they produce symmetric confidence intervals whose midpoints do not account for selection and therefore are biased; thus, the intervals must be wide enough to account for this bias. In this paper, we investigate infer-and-widen approaches in three vignettes: the winner’s curse, maximal contrasts, and inference after the lasso. In each of these examples, we show that a state-of-the-art infer-and-widen proposal leads to confidence intervals that are wider than a non-infer-and-widen alternative. Furthermore, even an “oracle” infer-and-widen confidence interval — the narrowest possible interval that could be theoretically attained via infer-and-widen — can be wider than the alternative.
infer-and-widen,
keywords:
[…]
, , , and
@@ 1 Introduction
To formalize this, consider a random variable $Y$ with domain $\mathcal{Y}$, and a collection of parameters $\theta$ whose elements can be indexed by a (possibly infinite) index set $\mathcal{I}$. For a fixed index $i\in\mathcal{I}$, and for $\alpha\in(0,1)$, a classical $1-\alpha$ confidence interval for $\theta_{i}$, which we will denote $\mathrm{CI}_{i}^{\alpha}$, satisfies
@@
Now, suppose that $i\in\mathcal{I}$ is no longer a fixed index; instead, some selection rule ${\mathcal{S}}:\mathcal{Y}\to\mathcal{I}$ uses the data to select a parameter on which to conduct inference. A confidence interval provides valid unconditional coverage [3] at level $\alpha$ if
@@
Unconditional coverage is not guaranteed by classical confidence intervals. In this paper, we consider methods that yield unconditional guarantees, in the sense of 1.2. Some of these confidence intervals also yield (stronger) conditional guarantees [10], in the sense that
$\mathbb{P}\left(\theta_{\mathcal{S}(Y)}\not\in\mathrm{CI}_{\mathcal{S}(Y)}^{\alpha}(Y)~|~\mathcal{C}(Y)\right)\leq\alpha$ (1.3)
for an event $\mathcal{C}(Y)$ satisfying $\sigma\left(\mathcal{C}(Y)\right)\supseteq\sigma\left({\mathcal{S}}(Y)\right)$, where $\sigma(\cdot)$ denotes the sigma algebra generated by a random variable, i.e., the coverage guarantee in 1.3 conditions on at least the selection event. By the law of total expectation, conditional coverage, 1.3, implies unconditional coverage, 1.2.
@@ 1.1 The infer-and-widen framework
As presented in Box 1.1, the infer-and-widen framework involves (Step 1) defining a target of inference using some or all of the data, and then (Step 2) symmetrically adjusting the critical values arising from “classical” inference to account for the fact that the target of inference is not fixed in order to achieve unconditional coverage, (1.2). The simplest examples of this strategy are the Bonferroni [4, 8], Holm-Bonferroni [14], and Šidák [23] procedures, which are suitable when a finite number of targets of inference are considered (and in the case of Šidák, under additional assumptions). In the special case of inference on contrasts of a linear model, the Scheffé correction is available [22]; this idea is further refined by the PoSI method [3]. While conceptually attractive, these approaches tend to be quite conservative (Bonferroni, Holm-Bonferroni, Šidák), are limited to a linear model setting (Scheffé, PoSI), or are often computationally intractable (PoSI).
Other recent infer-and-widen approaches have attempted to address these shortcomings. The local simultaneous inference approach of Zrnic and Fithian [26] focuses a simultaneous correction on the subset of hypotheses that are most likely given the observed data. The algorithmic stability proposal of [28] also adapts to the details of the selection algorithm. By drawing connections with the field of differential privacy [9], the “stability” of the approach for selecting the target of inference in Step 1 of Box 1.1 is quantified; this, in turn, leads to a stability-specific adjustment of the critical values in Step 2 of Box 1.1 (see also Section A of the online supplement). In a sense, these two approaches achieve the lofty goal laid out in the PoSI proposal of [3] by creating frameworks to compute selection-specific adjustments to the critical values.
Box 1: Infer-and-widen. 1. A target of inference is selected from the data. 2. “Classical” inference is conducted, with a symmetric adjustment of the critical values to account for the data-driven selection of the target of inference.
@@ 1.2 Other approaches for valid selective inference
Not all approaches that yield unconditional guarantees of the form (1.2) fall into the infer-and-widen framework. Some methods produce asymmetric confidence intervals around the biased point estimate: for instance, Benjamini et al. [2] and Fuentes et al. [11] construct an asymmetric Bonferroni correction that accounts for the direction of the bias. More recently, Zrnic and Fithian [27] extend the local simultaneous approach of Zrnic and Fithian [26] by additionally adapting the correction to the specific selection algorithm. The hybrid approaches of [1] and [18] apportion a fraction of the Type 1 error budget to rule out unlikely selection events, and then construct a conditionally-valid confidence interval on the likely hypotheses; note that this only guarantees unconditional coverage, 1.2. Lastly, there is a line of empirical Bayes work whose intervals are centered at shrinkage estimators which, while likely biased, improve upon the non-selective estimator [15].
Furthermore, a number of methods yield coverage guarantees conditional on the selection event, i.e., (1.3). Some methods do this by obtaining independent training and test sets; then, the training set can be used to define the target of inference, and the test set to conduct inference on this target. Because the two sets are independent, we can conduct inference using the test set as though the target of inference were fixed, despite the fact that it is actually a function of the training data. Sample splitting, applicable in the setting of independent and identically distributed observations, provides one approach to obtain independent training and test sets [6]. More recent data thinning proposals generalize the idea of sample splitting, and can be applied when only a single observation is available [21, 19, 7].
In cases where it is either impossible or undesirable to create independent training and test sets, strategies that achieve conditional guarantees of the form 1.3 may still be available. For instance, a target of inference can be defined using all of the data, and then inference can be conducted conditional on the event that the data led to this particular target of inference. Examples include the proposals of [10, 16, 5, 12, 1]. Alternatively, if a researcher is willing to define the target of inference using only a subset of the data or using a noisy version of the data, then data fission [17] or randomized conditional selective inference [24, 20] may be applicable. Randomization in this manner preserves information for inference, overcoming a limitation of conditioning on selection using all of the data.
@@ 1.3 Contributions and organization
In this paper, we conduct an empirical and theoretical evaluation of infer-and-widen strategies in the context of three vignettes. The first two vignettes involve inference after a randomized selection rule, and we derive alternative confidence intervals to compare to an existing infer-and-widen approach. In the third vignette, we compare a suite of existing methods. The bottom line of our paper is as follows: when comparing infer-and-widen strategies to alternative selective inference methods that make use of identical selection events, the latter often outperform the former in terms of confidence interval width. Furthermore, in some settings, the latter even outperform the “oracle” infer-and-widen procedure: the narrowest possible interval that could be theoretically attained within the infer-and-widen framework.
The organization is as follows. In Section 2, we formalize the infer-and-widen framework. In Sections 3, 4, and 5, we compare its performance to new alternative intervals that we derive in two vignettes considered in [28], and to existing alternative intervals in an example involving model selection via the lasso. In Sections 3 and 4, the selection event is randomized, whereas in Section 5 it is not. In Section 6, we show analytically that the infer-and-widen interval can suffer from a highly biased midpoint. We close with a discussion in Section 7. Theoretical results and additional empirical comparisons are proven in the online supplement. Scripts to fully reproduce all numerical results can be found at https://github.com/rflperry/infer_and_widen.
@@ 2 Formalizing the infer-and-widen framework
We now formalize the infer-and-widen (I&W) framework, outlined in Box 1.1. This involves a modification of the classical confidence interval $\mathrm{CI}_{i}^{\alpha}$ in (1.1) to account for the fact that $\theta_{\mathcal{S}(Y)}$ is a data-driven parameter. Based on the particulars of the selection function $\mathcal{S}(\cdot):\mathcal{Y}\rightarrow\mathcal{I}$ and data $Y$, a modified $\alpha$-level, $H(\alpha,{\mathcal{S}},Y)$, is derived so that $\mathrm{CI}^{H(\alpha,{\mathcal{S}},Y)}_{\mathcal{S}(Y)}(Y)$ is unconditionally valid at level $\alpha$ in the sense of 1.2, i.e.,
@@
In a Bonferroni correction we have $H(\alpha,{\mathcal{S}},Y)=\alpha/|\mathcal{I}|$, and in the Holm-Bonferroni correction it is slightly larger. Since these approaches adapt neither to the selection rule $\mathcal{S}(\cdot):\mathcal{Y}\rightarrow\mathcal{I}$ nor to the data $Y$, the constant $H(\alpha,{\mathcal{S}},Y)$ is easy to compute but overly conservative. Despite the notation $H(\alpha,{\mathcal{S}},Y)$ used in (2.1), the aforementioned infer-and-widen proposals result in confidence intervals whose widths do not depend on the particulars of the selection rule $\mathcal{S}(\cdot)$ nor data $Y$.
The foundational PoSI proposal [3] considers the setting of linear regression with fixed regressors. They derive a constant $H(\alpha,{\mathcal{S}},Y)$ that enables inference on the coefficient(s) in any submodel, regardless of the form of $\mathcal{S}(\cdot)$ or observed data $Y$. However, for a particular selection rule $\mathcal{S}(\cdot)$, the resulting confidence interval may be quite conservative, and computing the constant $H(\alpha,{\mathcal{S}},Y)$ is challenging or infeasible.
In contrast are two newer approaches, which provide narrower constants $H(\alpha,{\mathcal{S}},Y)$ that depend on the selection rule ${\mathcal{S}}(\cdot)$. The locally simultaneous inference method of Zrnic and Fithian [26] apportions the Type-1 error budget $\alpha$, first to use the data $Y$ to rule out the most unlikely hypotheses, and then to perform a simultaneous correction over a subset of all $\left|\mathcal{I}\right|$ hypotheses. The algorithmic stability proposal of [28] adjusts the critical value $H(\alpha,{\mathcal{S}},Y)$ based on how “stable” the selection rule ${\mathcal{S}}(\cdot)$ is; intuitively, a more “stable” selection rule requires a less substantial adjustment. Section A in the online supplement contains a more detailed overview.
@@ 3 Vignette #1: The winner’s curse
and the selection rule is
$\mathcal{S}(Y):=\mathop{\mathrm{arg\,max}}_{i\in[n]}Y_{i}.$ (3.2)
It is not hard to see that a $1-\alpha$ confidence interval for $\mu_{\mathcal{S}(Y)}$ that treats $\mathcal{S}(Y)$ as though it were fixed (rather than a function of the data) will fail to achieve the nominal coverage. This is the “winner’s curse” [1].
To achieve a “stable” algorithm, [28] consider a randomized version of the selection rule (3.2) via the addition of centered Laplace noise,
@@
Proposition 1 (An infer-and-widen interval based on (3.3), given by [28]).
@@
has $1-\alpha$ unconditional coverage for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ in the sense of 2.1.
In the context of Proposition 1, the selection algorithm in 3.3 is “stable” at level $\eta$ with probability at least $1-\nu$. We defer a more detailed discussion to Section A in the online supplement.
We will now derive an alternative interval that makes use of exactly the same selection event, (3.3), using ideas from the data fission proposal of [17]. Towards this end, note that $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)=\mathcal{S}(Y^{\mathrm{train}})$ where $Y^{\mathrm{train}}:=Y+\zeta$. Since $Y\mid Y^{\mathrm{train}}$ does not admit a tractable distribution, we will additionally condition on $\left({\rm sign}(\zeta_{1}),\ldots,{\rm sign}(\zeta_{n})\right)$. This yields a multivariate truncated normal distribution with independent marginals.
@@
has $1-\alpha$ conditional coverage for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ in the sense of 1.3 with $\mathcal{C}(Y)=\left(Y+\zeta,\delta\right).$
@@
Thus, both (3.4) and (3.7) provide confidence intervals for $\mu_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ that attain $1-\alpha$ unconditional coverage. We should prefer the approach that yields narrower intervals. Figure 1(a) displays the ratio of the widths of the two intervals when $\mu=0$ and $\sigma^{2}=1$, as we vary the sample size $n$ and the variance $2c^{2}$ of the Laplace noise in (3.3).
The infer-and-widen interval (3.4) involves tuning parameters $\nu$ and $\eta$; to construct Figure 1(a), for a fixed value of the noise $c$ and sample size $n$ in (3.3), we select the $\nu$ and $\eta$ that minimize the confidence interval width, subject to the constraints in Proposition 1. The data fission intervals (3.7) are much narrower than the infer-and-widen intervals (3.4), across all values of $n$ and $c$ considered.
It is possible that the width of the infer-and-widen interval (3.4) is due to the fact that [28] apply a series of inequalities in its derivation; furthermore, it is natural to wonder whether another infer-and-widen proposal could perform better. Therefore, we now consider the “oracle” infer-and-widen interval: this is the narrowest possible interval centered at $Y_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}$ that achieves a given coverage. It is not computable in practice, as it requires knowledge of the true parameter value $\mu$. In Figures 1(b, c), we numerically compare the “oracle” infer-and-widen method to the data fission interval (3.7) under the selection rule (3.3) with variances $2c^{2}=0.33$ and $3$, respectively. The former method has a wider average width than the latter: thus, no infer-and-widen interval can outperform the data fission interval (3.4) in this setting.
Section C of the supplementary materials contains a related comparison of methods for inference after non-randomized selection using 3.2. The conclusions are the same.
Figure 1: Vignette #1 under (3.1) with $\mu=0$, $\sigma=1$, and the selection rule in $3.$ Results are averaged over (a) $250$ or (b, c) $1000$ simulated datasets. (a): We display the ratio of the average widths of the infer-and-widen (I&W) 3.4 and data fission (DF) 3.7 intervals with $95\%$ coverage across values of the sample size $n$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the infer-and-widen interval is wider. (b, c): Fixing $n=100$, we display the average width of the “oracle” infer-and-widen interval and the data fission interval 3.7 as a function of coverage. The 95% classical interval and the “attainable” 95% infer-and-widen interval based on algorithmic stability 3.4 are also displayed. The Laplacian noise variance is low and high in panels (b) and (c), respectively.